The freezing point depression constant for water is 1.86 oC/mol of particles. If 345 g of calcium chloride (CaCl2) salt is dissolved in 1000 g of water, what will be the freezing pointof the solution?


1
Expert's answer
2021-04-02T05:17:20-0400

"\Delta"tf=ikfm


Kf=1.86


i= 1+(3-1)"\alpha" .......... {"\alpha" =1}

"\implies" i = 3


m = "n_{solute\ufeff}\\over {m_{H_2O(kg)}}" "\over"


m="3.1\over 1 (kg)" ............{nsolute= "345\over 111" =3.1}


"\implies" m= 3.1


Δtf="3\times1.86\times3.1"


=17.298

Answer = 0-17.298

= -17.298